write in standard form,an ellipse with foci (-3,2) and (1,2) and major axis of length 8
Answer by lwsshak3(11628)

write in standard form,an ellipse with foci (-3,2) and (1,2) and major axis of length 8
**
Since x-coordinates of the foci change but y-coordinates do not, major axis is horizontal.
Standard form of equation for an ellipse with horizontal major axis: (x-h)^2/a^2+(y-k)^2/b^2=1
x-coordinate of center is half-way between x-coordinates of foci=-1
y-coordinate of center is y-coordinate of foci=2
center: (-1,2)
length of major axis=8=2a
a=4
a^2=16
..
distance between x-coordinates of foci=4=2c
c=2
c^2=4
..
c^2=a^2-b^2
b^2=a^2-c^2=16-4=12
..
Equation of given ellipse:
(x-h)^2/a^2+(y-k)^2/b^2=1
(x+1)^2/16+(y-2)^2/12=1